OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 4 points) What is four divided by two-fifths? The model shows four whole strips lying across. Each whole strip is labeled as one. Below each whole strip are five one-fifth strips lying across. Each one-fifth strip is labeled as one-fifth. The one-fifth strips are also grouped by bold outlines in twos. 5 8 ten twenty

OpenStudy (anonymous):

ten I believe

OpenStudy (anonymous):

your not suppose to give out answers

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

4/(2/5)=4×(5/2)=10

OpenStudy (anonymous):

:D